OpenStudy (anonymous):

the average weight in pounds for a man can be estimated by the following function where x represents the mans height in inches and f(x) is his weight in pounds. f(x)=0.117x^1.7 a. what is the average weight of a 68 inch tall man?? b. what is teh height of a man that weighs 185 pounds??

OpenStudy (anonymous):

a)0.177*(68)^1.7 b)(185/0.117)^(1/1.7)=x

OpenStudy (anonymous):

\[f[x] \text{=} 0.117*x{}^{\wedge}1.7 \] Rewrite the equation with rationalized decimal numbers and simplify \[f[x]==\frac{117 x^{17/10}}{1000} \] \[f[68]==\frac{1989}{125} 2^{2/5} 17^{7/10}=152.6 \] \[\text{Solve}\left[\frac{117 x^{17/10}}{1000}=185,x\right], x=\frac{50\ 2^{13/17} 5^{6/17} 37^{10/17}}{3\ 3^{3/17} 13^{10/17}}=76.2\] Solved with Mathematica 8.

OpenStudy (anonymous):

\[f[x] \text{=} 0.117*x{}^{\wedge}1.7 \] Rewrite the equation with rationalized decimal numbers and simplify \[f[x]==\frac{117 x^{17/10}}{1000} \] \[f[68]==\frac{1989}{125} 2^{2/5} 17^{7/10}=152.6 \] \[\text{Solve}\left[\frac{117 x^{17/10}}{1000}=185,x\right], x=\frac{50\ 2^{13/17} 5^{6/17} 37^{10/17}}{3\ 3^{3/17} 13^{10/17}}=76.2\] Solved with Mathematica 8.

OpenStudy (anonymous):

Sorry for the double posting. Completely unintentional.

OpenStudy (anonymous):

its fine. thanks for help

OpenStudy (anonymous):

You are welcome.